The cost of 8 CD's plus a sales tax of $1.34 is $105.26. What equation can I use to solve this?
Answer by flaminhart(1)

8x+1.34+105.26
subtract 1.34 from both sides. 8x=105.26-1.34
8x=103.92
divide each side by 8
8x/8=103.92/8
the price is 12.99 each